OpenStudy (anonymous):

Which of the following is a solution of y > |x| - 6? (-5, 1) (-1, -5) (5, -1) b

OpenStudy (anonymous):

A

OpenStudy (anonymous):

umumum

OpenStudy (anonymous):

me?

OpenStudy (nincompoop):

show your solutions or I'll have you banned for giving out answers

OpenStudy (anonymous):

or him

OpenStudy (anonymous):

me or him...im scared now

OpenStudy (anonymous):

i got b when i worked it ut i was just asking if it was correct

terenzreignz (terenzreignz):

All the same... how DID you get b?

OpenStudy (anonymous):

1-6 =-5 rite ????

OpenStudy (anonymous):

is the answer given wrong????

OpenStudy (anonymous):

thats what my calc

OpenStudy (nincompoop):

no one is allowed to give out answers @vivek4607 coocoobird

OpenStudy (luigi0210):

@vivek4607 no giving away answers, like nin stated

terenzreignz (terenzreignz):

... it would be useful to point out that you, @dirtydan667 , are not currently dealing with an equation, but rather, an inequality, and a STRICT inequality, at that (>, not \(\ge\))

OpenStudy (anonymous):

it would be x-6<y, and -6-x>y (i think :P)

OpenStudy (anonymous):

um um um um um um um um im scured

OpenStudy (anonymous):

@nincompoop ...dude...she just asked for the solution...not for the how solution comes....better read the question first....:)

OpenStudy (anonymous):

who she

terenzreignz (terenzreignz):

Solution = Answer... and...

OpenStudy (anonymous):

ohk...thanx for reminding me...:)

OpenStudy (anonymous):

wait im confused what is the awnsere

terenzreignz (terenzreignz):

@dirtydan667 Listen up... to know if a point is a solution of an equation (or an inequality) just plug everything in and see if it all checks out. Here's an example... try the point \((\color{red}3,\color{blue}6) \) \[\Large \color{blue}y > |\color{red}x|- 6\]Plug in... \[\Large \color{blue}6 > |\color{red}3| - 6\] Simplify... \[\Large 6 > -3\] Which is true... so, the point \((\color{red}3,\color{blue}6) \) IS a solution. Now try with your actual choices.

OpenStudy (anonymous):

y > |x| - 6 y+6>|X| y+6=x OR y+6=-x y+6=x or -y-6=x by putting all the values of x and y given u can find out ur answer

OpenStudy (anonymous):

so A ?

OpenStudy (anonymous):

u can check it....:)

terenzreignz (terenzreignz):

so B? or maybe even C? Test it.

OpenStudy (nincompoop):

try to SOLVE it

OpenStudy (nincompoop):

you don't want to be banned for soliciting answers now

OpenStudy (anonymous):

so it is a 1>-1

terenzreignz (terenzreignz):

Where did that come from? D:

OpenStudy (anonymous):

i enterd the points from A

terenzreignz (terenzreignz):

Show.

OpenStudy (anonymous):

kk one sec

OpenStudy (nincompoop):

eyes are unto you, coocoobird, you better show your solution or just close this post and move on

OpenStudy (anonymous):

um ok first i enterd the points in a then did the || then i got 1>-1

OpenStudy (anonymous):

whats so hard to beileve

OpenStudy (nincompoop):

eh? that's not a solution :) if you've read terenz' example, then you would do the same format for your solution. c'mon, man, you can do this…. this question is beneath your ability

OpenStudy (anonymous):

no its not the soultuion but it means a is correct

OpenStudy (nincompoop):

I wouldn't know until you start plugging in values

OpenStudy (debbieg):

No @vivek, y+6>|X| is NOT equivalent to y+6=x OR y+6=-x An abs value inequality does not become a compound equality. think about what is means that: y+6>|X| It means that the DISTANCE of X from 0 is less than y+6. Therefore: -(y+6)<x<(y+6) or as two separate inequalities you have: -y-6<x and x<y+6 Now if you didn't simply have points to check, you would solve each, and take the INTERSECTION as your solution set. This can be illustrated graphically by graphing the 2 lines in the x-y plane and shading in the area where all the points make both inequalities true.